# On Scaling LT-Coded Blockchains in Heterogeneous Networks and their Vulnerabilities to DoS Threats

### v2

@@ 3.4 Complexity vs Overhead
In Fig. 5, we pictorially depict the trade-off between the computational complexity and bootstrap overhead of the various decoders when $k=500$ and when their decoding failure rate is upper bounded by $\delta=0.1$. In this depiction, the blue circle in the extreme right corner represents the BP decoder which clearly has the lowest computational complexity, however, a very high bootstrap overhead. It is vice-versa for the OFG decoder, which is represented using the red square. Labels in black and magenta denote the overhead-complexity pairs of the family of BRH decoders with their $K_{BR}$ values ranging from $K_{OFG}$ to $K_{BP}$, and the family of CRH decoders with their $\eta_{c}$ values ranging from 0 to $k$, respectively. It can be observed that the respective families of the BRH and CRH decoders act as bridges between the OFG and BP decoders, in terms of the overhead and complexity trade-off. Therefore, depending on whether the bucket node wants to operate at a fixed bootstrap overhead or fixed computational complexity, it can choose between the BRH and CRH decoders. If the bucket node chooses BRH decoder, it needs to select the parameter $K_{BR}$ that allows it to achieve its desired combination of computational complexity and bootstrap overhead. On the other hand, if CRH decoder is chosen, the bucket node needs to select the parameter $\eta_{c}$ that satisfies its complexity and overhead requirements.
@@ 3.5 The Mirroring Cost
A bucket node is always associated with a feasible range of computational complexity and bootstrap overhead over which it could operate. Out of the multiple complexity-overhead pairs which falls within this feasible range, a bucket node has to select one pair which gives the optimal performance. This could be done only after taking into consideration the costs associated with the bucket node in performing computations as well as downloading data. In this line, we define a new metric known as the absolute mirroring cost, which encompasses the costs associated with computation and download for the bucket node, along with the computational complexity and communication overhead of the decoder. Henceforth, we use the absolute mirroring cost as the standard metric to compare the performances of the LT decoders. To quantify this metric, we represent $c_{1}$ as the cost associated with the bucket node performing one XOR operation between two droplets. Similarly, $c_{2}$ represents the cost for the bucket node to contact a droplet node and download one droplet from it. With that, the absolute mirroring cost is defined as follows:
@@
In the rest of this section, we analyse the absolute mirroring cost using the cost ratio $\alpha$, where $\alpha=\frac{c_{2}}{c_{1}}$. In particular, we normalise $c_{1}$ to 1 and evaluate the normalized absolute mirroring costs of the decoders as functions of $\alpha$. We will henceforth refer to the normalized absolute mirroring cost as “mirroring cost”. From Definition 3, the mirroring costs of BP and OFG decoders denoted by $M_{BP}(\alpha)$ and $M_{OFG}(\alpha)$ respectively, are:
@@ 3.5.1 Mirroring Costs of BRH and CRH Decoders
The mirroring costs of BRH and CRH decoders, indicated by $M_{BR}(\alpha)$ and $M_{CR}(\alpha)$, respectively, are given in (5) and (6), respectively.
$M_{BR}(\alpha)=K_{BR}\cdot D\cdot\frac{\operatorname{\mathbb{E}}[\eta_{B}]}{k}+\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]+\alpha\cdot K_{BR}$ (5)
$M_{CR}(\alpha)=K_{CR}^{av}\cdot D\cdot\frac{\eta_{c}}{k}+(k-\eta_{c})^{2}+\alpha\cdot K_{CR}^{av}$ (6)
It is important to note that the mirroring costs of BRH and CRH decoders are not fixed for a given $\alpha$, unlike the BP and OFG decoders, which have static mirroring costs. This is due to the BRH decoder’s variable $K_{BR}$, which can range from $K_{OFG}$ to $K_{BP}$, and CRH decoder’s variable $\eta_{c}$ which can range from 00 to $k$. Therefore, we propose an optimization problem (Problem 1) for the BRH decoder to obtain the optimal number of droplet nodes that must be contacted by the bucket node, denoted by $K_{min}(\alpha)$, in order to achieve the minimum mirroring cost of the decoder. Similarly, we also propose an optimization problem (Problem 2) for the CRH decoder to obtain the optimal $\eta_{c}$, denoted by $\eta_{min}(\alpha)$, in order to achieve the minimum mirroring cost of the decoder.
{mdframed}
@@
For a given $\alpha>0$, solve
$K_{min}(\alpha)=$ $\,\underset{K_{BR}}{\arg\min}\left\{K_{BR}D\frac{\operatorname{\mathbb{E}}[\eta_{B}]}{k}+\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]+\alpha K_{BR}\right\}$
s.t. $K_{OFG}\leq K_{BR}\leq K_{BP}.\;$
{mdframed}
@@
For a given $\alpha>0$, solve
$\eta_{min}(\alpha)=$ $\,\underset{\eta_{c}}{\arg\min}\left\{K_{CR}^{av}\cdot D\cdot\frac{\eta_{c}}{k}+(k-\eta_{c})^{2}+\alpha\cdot K_{CR}^{av}\right\}$
s.t. $0\leq\eta_{c}\leq k.\;$
@@ 4 Denial-of-Service Threats on LT Coded Blockchains
Despite the fact that LT coded blockchains facilitate seamless scalability, DoS threats pose a critical risk to this feature, disrupting their ability to operate effectively in decentralized environments. In this respect, we consider an adversary or a group of adversaries that launch DoS attacks [21, 22, 23, 24, 25] on a subset of droplet nodes, thereby preventing new bucket nodes from using the compromised droplet nodes for their decoding process. As a consequence, blockchain scalability is hindered, as new bucket nodes who want to join the network may not receive sufficient data to retrieve the original blockchain.
Towards highlighting the vulnerability of coded blockchains to DoS threats, one can classify the threat model that an adversary undertakes into two broad categories based on how he selects a subset of droplet nodes to launch the attack: (i) Oblivious model and (ii) Non-oblivious model. In this context, an adversary employing the oblivious model, referred to as an oblivious adversary does not observe the storage contents of any of the droplet nodes before performing the attack. Consequently, the oblivious adversary blindly launches DoS attack on a random subset of droplet nodes. Henceforth, we refer to an attack executed by an oblivious adversary as blind attack. The authors of [17] have shown that LT coded blockchains are resilient against such blind attacks. However, to the best of our knowledge, the impact of a non-oblivious adversarial model on LT-coded blockchains has yet to be investigated.
@@
Pointing at this research gap, in this work, we adopt a non-oblivious model, wherein the adversary gains access to a set of droplet nodes, reads their contents, and uses this information to select a subset of the accessed nodes to attack. Henceforth, we refer to the attacked droplet nodes as erased nodes, and the residual droplet nodes that are not attacked as honest nodes. In the subsequent section, we show that the selection of the subset of droplet nodes to be attacked is determined by the decoder employed by the bucket node.
@@ 4.1 DoS Threats from Non-Oblivious Adversaries
[…]
@@ 4.1.1 Attack Strategy against BP Decoder
[…]
BP_Optimal_Attack A reasonable attack strategy against the BP decoder is to arrange the $\sigma_{0}S$ droplet nodes in the increasing order of their degrees and then erase the first $\sigma S$ droplet nodes of the sorted set. Note that the order of the droplet nodes having the same degree can be chosen in an arbitrary manner.
@@
[…]
@@
[…]
BP_Score_Attack The score based attack strategy against the BP decoder involves the following steps.
@@
      For each droplet node $\mathbf{y}\in\mathcal{X}_{d_{i}}$ with degree $\tilde{d}$, if the neighbours of $\mathbf{x}_{i}$ are a subset of the neighbours of $\mathbf{y}$, the adversary updates the score of $\mathbf{x}_{i}$ as $\mathcal{S}(\mathbf{x}_{i})$leftarrow $\mathcal{S}(\mathbf{x}_{i})$ + $\left(\frac{1}{2}\right)^{(\tilde{d}-1)}$. The adversary repeats this step for all $\mathbf{y}\in\mathcal{X}_{d_{i}}$.
@@
Block 1 00 00 $1$ 00 00 00 00 $1$ $1$ $1$
Block 2 00 00 00 00 00 $1$ $1$ 00 $1$ $1$
Block 3 $1$ 00 $1$ 00 00 00 $1$ $1$ $1$ $1$
Block 4 00 00 00 $1$ 00 00 00 $1$ 00 $1$
Block 5 00 $1$ 00 00 $1$ $1$ 00 00 $1$ $1$
Block 6 00 00 00 $1$ $1$ 00 $1$ $1$ 00 $1$
@@
In the first step, the adversary calculates the initial scores of droplet nodes $x_{i}$ for $i\in[1,10]$, as $\left(\frac{1}{2}\right)^{(d-1)}$, where $d$ is the degree of the $i^{th}$ droplet node. These initial scores are presented in Table 2. In the next step, the adversary starts with $\mathbf{x}_{1}$ to perform score updation. Since the degree of $\mathbf{x}_{1}$ is $1$, $\mathcal{X}_{1}=\{\mathbf{x}_{i}~{}|~{}i\in[3,10]\}$. We observe that the neighbours of $\mathbf{x}_{1}$ are a subset of the neighbours of $\mathbf{x}_{3}$, $\mathbf{x}_{7}$, $\mathbf{x}_{8}$, $\mathbf{x}_{9}$ and $\mathbf{x}_{10}$ of the set $\mathcal{X}_{1}$, whose degrees are $2$, $3$, $4$, $4$ and $6$ respectively. As a result, the updated score of $\mathbf{x}_{1}$ is $1+(\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{8}+\frac{1}{32})$. Similarly, the updated scores of the remaining droplet nodes, i.e. $\mathbf{x}_{i}$ for $i\in[2,10]$ are also calculated. The final scores of droplet nodes read by the adversary, $\mathbf{x}_{i}$ for $i\in[1,10]$ are {$2.03125$, $2.15625$, $0.78125$, $0.65625$, $0.53125$, $0.65625$, $0.28125$, $0.15625$, $0.15625$, $0.03125$}, respectively. Finally, the adversary arranges the droplet nodes in the decreasing order of their respective scores to obtain the new order as $\mathbf{x}_{2}$, $\mathbf{x}_{1}$, $\mathbf{x}_{3}$, $\mathbf{x}_{4}$, $\mathbf{x}_{6}$, $\mathbf{x}_{5}$, $\mathbf{x}_{7}$, $\mathbf{x}_{8}$, $\mathbf{x}_{9}$, $\mathbf{x}_{10}$}. He then erases the first $5$ droplet nodes of this ordered set.
In Section 4.2, we will show that the score based attack strategy performs better than the degree based attack strategy in terms of enhancing the failure rate at a legitimate bucket node.

### v3

@@ 3.4 Complexity vs Overhead
In Fig. 5, we pictorially depict the trade-off between the computational complexity and bootstrap overhead of the various decoders when $k=500$ and when their decoding failure rate is upper bounded by $\delta=0.1$ along with the RSD parameter c=1c=1.111Note that the RSD parameters are fixed to $\delta=0.1$ and $c=1$ for the presentation of results. Since the behavior of the bootstrap overhead and the average computational complexity of the decoders are already well established, similar experiments can be conducted for the other values of $\delta$ and $c$. In this depiction, the blue circle in the extreme right corner represents the BP decoder which clearly has the lowest computational complexity, however, a very high bootstrap overhead. It is vice-versa for the OFG decoder, which is represented using the red square. Labels in black and magenta denote the overhead-complexity pairs of the family of BRH decoders with their $K_{BR}$ values ranging from $K_{OFG}$ to $K_{BP}$, and the family of CRH decoders with their $\eta_{c}$ values ranging from 0 to $k$, respectively. It can be observed that the respective families of the BRH and CRH decoders act as bridges between the OFG and BP decoders, in terms of the overhead and complexity trade-off. Therefore, depending on whether the bucket node wants to operate at a fixed bootstrap overhead or fixed computational complexity, it can choose between the BRH and CRH decoders. If the bucket node chooses BRH decoder, it needs to select the parameter $K_{BR}$ that allows it to achieve its desired combination of computational complexity and bootstrap overhead. On the other hand, if CRH decoder is chosen, the bucket node needs to select the parameter $\eta_{c}$ that satisfies its complexity and overhead requirements.
@@ 3.5 The Mirroring Cost
A bucket node is always associated with a feasible range of computational complexity and bootstrap overhead over which it could operate. Out of the multiple complexity-overhead pairs which falls within this feasible range, a bucket node has to select one pair which gives the optimal performance. This could be done only after taking into consideration the costs associated with the bucket node in performing computations as well as downloading data. In this line, we define a new metric known as the absolute mirroring cost, which encompasses the costs associated with computation and download for the bucket node, along with the computational complexity and communication overhead of the decoder. Henceforth, we use the absolute mirroring cost as the standard metric to compare the performances of the LT decoders. To quantify this metric, we represent $c_{1}>0$ as the cost associated with the bucket node performing one XOR operation between two droplets. Similarly, $c_{2}>0$ represents the cost for the bucket node to contact a droplet node and download one droplet from it. With that, the absolute mirroring cost is defined as follows:
@@
In the rest of this section, we analyse the absolute mirroring cost using the cost ratio $\alpha$, where $\alpha=\frac{c_{2}}{c_{1}}$. In particular, we normalise $c_{1}$ to 1 and evaluate the normalized absolute mirroring costs of the decoders as functions of $\alpha$. We will henceforth refer to the normalized absolute mirroring cost as “mirroring cost”. From Definition 3, for a given $\alpha>0$, the mirroring costs of BP and OFG decoders denoted by $M_{BP}(\alpha)$ and $M_{OFG}(\alpha)$ respectively, are:
@@ 3.5.1 Mirroring Costs of BRH and CRH Decoders
For a given $\alpha>0$, the mirroring costs of BRH and CRH decoders indicated by $M_{BR}(\alpha)$ and $M_{CR}(\alpha)$, respectively, are given in (13) and (14), respectively.
$M_{BR}(\alpha)=K_{BR}\cdot D\cdot\frac{\operatorname{\mathbb{E}}[\eta_{B}]}{k}+\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]+\alpha\cdot K_{BR}.$ (13)
$M_{CR}(\alpha)=K_{CR}^{av}\cdot D\cdot\frac{\eta_{c}}{k}+(k-\eta_{c})^{2}+\alpha\cdot K_{CR}^{av}.$ (14)
It is important to note that the mirroring costs of BRH and CRH decoders are not fixed for a given $\alpha$, unlike the BP and OFG decoders, which have static mirroring costs. This is due to the BRH decoder’s variable $K_{BR}$, which can range from $K_{OFG}$ to $K_{BP}$, and CRH decoder’s variable $\eta_{c}$ which can range from 0 to $k$. Therefore, we propose an optimization problem (Problem 1) for the BRH decoder to obtain the optimal number of droplet nodes that must be contacted by the bucket node, denoted by $K_{min}(\alpha)$, in order to achieve the minimum mirroring cost of the decoder. Similarly, we also propose an optimization problem (Problem 2) for the CRH decoder to obtain the optimal $\eta_{c}$, denoted by $\eta_{min}(\alpha)$, in order to achieve the minimum mirroring cost of the decoder.
@@
For a given $\alpha>0$, solve $K_{min}(\alpha)=$ $\,\underset{K_{BR}}{\arg\min}\left\{K_{BR}D\frac{\operatorname{\mathbb{E}}[\eta_{B}]}{k}+\operatorname{\mathbb{E}}[(k-\eta_{B})^{2}]+\alpha K_{BR}\right\}$ s.t. $K_{OFG}\leq K_{BR}\leq K_{BP}.\;$
@@
For a given $\alpha>0$, solve $\eta_{min}(\alpha)=$ $\,\underset{\eta_{c}}{\arg\min}\left\{K_{CR}^{av}\cdot D\cdot\frac{\eta_{c}}{k}+(k-\eta_{c})^{2}+\alpha\cdot K_{CR}^{av}\right\}$ s.t. $0\leq\eta_{c}\leq k.\;$
@@ 4 Denial-of-Service Threats on LT Coded Blockchains
In the preceding sections, we have shown that LT coded blockchains support low-complexity algorithms for the bucket nodes to download and mirror the transactions of the blockchains. As a consequence, bucket nodes, despite possessing limited resources in terms of computational workload and bandwidth, can join the blockchain thereby facilitating seamless scalability. Despite these benefits of LT coded blockchains, they may also be vulnerable to a wide range of security threats, akin to any other networked systems. In the context of blockchains, some relevant security threats could be those degrading the infrastructure scalability or compromising the ability to perform legitimate transactions.
In this work, we are interested in studying the vulnerabilities of LT coded blockchains against DoS (denial of service) threats that pose a critical risk to the scalability of LT coded blockchains, disrupting their ability to operate effectively in decentralized environments. In particular, our vulnerability analysis is on scenarios where the attacker’s objective is to design strategies that render a subset of the droplet nodes unavailable for other bucket nodes to download the blockchain. As a consequence, new nodes will not be able to join the blockchain. Thus, the number of nodes that store the complete history of transactions in the blockchain will not grow thereby limiting the number of honest nodes to achieve consensus for verifying the transactions. Although the direct consequence of the attacker’s objective is to limit the scalability of LT coded blockchain, its indirect and long-term impact is to curtail the network’s ability for consensus due to the presence of fewer number of full nodes.
In the following section, we present the assumptions we make with respect to the capabilities of the attacker, in terms of resources to execute the attacks, and the implementation methods through which the attacks could be executed in practice.
4.1 Threat Model
We consider an adversary or a group of adversaries that launch DoS attacks [21, 22, 23, 24, 25] on a subset of droplet nodes, thereby preventing new bucket nodes from using the compromised droplet nodes for their decoding process. As a consequence, blockchain scalability is hindered, as new bucket nodes who want to join the network may not receive sufficient data to retrieve the original blockchain. To launch DoS threats, an adversary may flood a targeted set of droplet nodes with repeated service requests, such that a legitimate bucket node is not able to access the compromised droplet nodes [21]. As an alternative, an adversary could also make the droplet nodes unavailable by compromising them akin to the methods used in ransomware attacks. We remark that the way an adversary makes the droplet nodes unavailable is out of the scope of this work. Instead, we argue that the way in which an adversary chooses the target droplet nodes to compromise is more relevant. This is because if the choice of the target droplet nodes is made based on the contents stored in them, then it might have a higher impact.
Towards choosing the target droplet nodes, one can classify the threat model into two categories, namely: (i) Oblivious model and (ii) Non-oblivious model. An adversary employing the oblivious model, referred to as an oblivious adversary, does not observe the storage contents of any of the droplet nodes before performing the attack. Consequently, the oblivious adversary blindly launches DoS attack on a random subset of droplet nodes. Henceforth, we refer to an attack executed by an oblivious adversary as blind attack. On the other hand, non-oblivious threat model refers to a scenario wherein an adversary gains access to a set of droplet nodes, reads their contents, and uses this information to select a subset of the accessed nodes to attack. In practice, a typical way for the adversary to gain access into the droplet nodes is to enter the network as a potential bucket node and then download their contents. After reading the contents of the droplet nodes, it could then decide to attack a subset of them.
@@ 4.1 Threat Model
The authors of [17] have shown that LT coded blockchains are resilient against such blind attacks. However, to the best of our knowledge, the impact of a non-oblivious adversarial model on LT-coded blockchains has yet to be investigated. Pointing at this research gap, in this work, we adopt a non-oblivious model, wherein the adversary gains access to a set of droplet nodes, reads their contents, and uses this information to select a subset of the accessed nodes to attack. To realize a non-oblivious model, we assume that the adversary in our threat model has sufficient resources to gain access into a set of droplet nodes, and read their contents. In practice, when the adversary joins the network as a potential bucket node, the resources required for this task corresponds to the communication-overheads for I/O access and download. Subsequently, based on the contents in the droplets, we assume that the adversary has necessary additional resources required to make the droplets unavailable, either through ransomware- or flooding-based attacks. Given that the resources for accessing the contents and executing attacks are attacker dependent, we parameterize such capabilities, and then study their impact on LT coded blockchains. Henceforth, we refer to the attacked droplet nodes as erased nodes, and the residual droplet nodes that are not attacked as honest nodes. In the subsequent section, we show that the selection of the subset of droplet nodes to be attacked is determined by the decoder employed by the bucket node.
@@ 4.2 DoS Threats from Non-Oblivious Adversaries
Let us consider a full node that contains $S$ number of droplet nodes under the LT coded architecture. Recall from Section 2 that the degree information of droplet nodes is available in the form of a binary vector $\mathbf{v}$, which is stored along with the coded blocks. On this full node, we consider a non-oblivious adversary who acts as a legitimate bucket node and reads the degree information of a randomly chosen $\sigma_{0}$-fraction of $S$ droplet nodes, where $\sigma_{0}\in(0,1)$. He then uses this information to tactically choose a subset of $\sigma S$ droplet nodes to erase among them, where $0<\sigma\leq\sigma_{0}$. Fig. 9 depicts our adversarial model wherein regions I, II and III respectively represent the set of droplet nodes that the adversary has (i) neither read nor erased, (ii) read, however not erased, and (iii) both read and erased. As a special case, note that our adversarial model will collapse to the oblivious model if $\sigma_{0}=\sigma$. This is equivalent to the adversary attacking a randomly chosen $\sigma$-fraction of $S$ droplet nodes without reading the degree information on their coded structures.
@@ 4.2.1 Attack Strategy against BP Decoder
A reasonable attack strategy against the BP decoder is to arrange the $\sigma_{0}S$ droplet nodes in the increasing order of their degrees and then erase the first $\sigma S$ droplet nodes of the sorted set. Note that the order of the droplet nodes having the same degree can be chosen in an arbitrary manner.
@@
Although the aforementioned degree based attack strategy effectively exploits the weakness of the BP decoder, a drawback of this strategy is that it considers droplet nodes with the same degree to be equally valuable to the BP decoder. However, it does not take into account the information about the neighbours of the droplet nodes. Recall that neighbours of a droplet node refer to the collection of input blocks of an epoch that were XORed to obtain that droplet node. In particular, if an adversary needs to erase a subset of droplet nodes, all having the same degree $d$ for some $d\in[k]$, ?THM? LABEL:BP_Optimal_Attack instructs him to erase a subset of them chosen at random. However, the droplet nodes that are not erased by the adversary may be more valuable to the BP decoder than those that are erased, thereby benefiting the legitimate bucket nodes.
@@
The score based attack strategy against the BP decoder involves the following steps.
@@
      For each droplet node $\mathbf{y}\in\mathcal{X}_{d_{i}}$ with degree $\tilde{d}$, if the neighbours of $\mathbf{x}_{i}$ are a subset of the neighbours of $\mathbf{y}$, the adversary updates the score of $\mathbf{x}_{i}$ as $\mathcal{S}(\mathbf{x}_{i})$ $\leftarrow$ $\mathcal{S}(\mathbf{x}_{i})$ + $\left(\frac{1}{2}\right)^{(\tilde{d}-1)}$. The adversary repeats this step for all $\mathbf{y}\in\mathcal{X}_{d_{i}}$.
@@
Block 1 0 0 11 0 0 0 0 11 11 11
Block 2 0 0 0 0 0 11 11 0 11 11
Block 3 11 0 11 0 0 0 11 11 11 11
Block 4 0 0 0 11 0 0 0 11 0 11
Block 5 0 11 0 0 11 11 0 0 11 11
Block 6 0 0 0 11 11 0 11 11 0 11
@@
In the first step, the adversary calculates the initial scores of droplet nodes $x_{i}$ for $i\in[1,10]$, as $\left(\frac{1}{2}\right)^{(d-1)}$, where $d$ is the degree of the $i^{th}$ droplet node. These initial scores are presented in Table 2. In the next step, the adversary starts with $\mathbf{x}_{1}$ to perform score updation. Since the degree of $\mathbf{x}_{1}$ is $1$, $\mathcal{X}_{1}=\{\mathbf{x}_{i}~|~i\in[3,10]\}$. We observe that the neighbours of $\mathbf{x}_{1}$ are a subset of the neighbours of $\mathbf{x}_{3}$, $\mathbf{x}_{7}$, $\mathbf{x}_{8}$, $\mathbf{x}_{9}$ and $\mathbf{x}_{10}$ of the set $\mathcal{X}_{1}$, whose degrees are $2$, 33, 44, 44 and $6$ respectively. As a result, the updated score of $\mathbf{x}_{1}$ is $1+(\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{8}+\frac{1}{32})$. Similarly, the updated scores of the remaining droplet nodes, i.e. $\mathbf{x}_{i}$ for $i\in[2,10]$ are also calculated. The final scores of droplet nodes read by the adversary, $\mathbf{x}_{i}$ for $i\in[1,10]$ are {$2.03125$, $2.15625$, $0.78125$, $0.65625$, $0.53125$, $0.65625$, $0.28125$, $0.15625$, $0.15625$, 0.031250.03125}, respectively. Finally, the adversary arranges the droplet nodes in the decreasing order of their respective scores to obtain the new order as $\mathbf{x}_{2}$, $\mathbf{x}_{1}$, $\mathbf{x}_{3}$, $\mathbf{x}_{4}$, $\mathbf{x}_{6}$, $\mathbf{x}_{5}$, $\mathbf{x}_{7}$, $\mathbf{x}_{8}$, $\mathbf{x}_{9}$, $\mathbf{x}_{10}$}. He then erases the first $5$ droplet nodes of this ordered set.
In Section 4.3.2, we will show that the score based attack strategy performs better than the degree based attack strategy in terms of enhancing the failure rate at a legitimate bucket node.